Determine the slope of the line containing the points (6,-2) and (-1,5).
Determine an equation for a line with slope 1/2 and y-intercept at (0, -3).

Answer by stanbon(57361)

Determine the slope of the line containing the points (6,-2) and (-1,5).
slope = [5--2]/[-1-6] = 7/-7 = -1
------------------------------
Determine an equation for a line with slope 1/2 and y-intercept at (0, -3).
Form of the line is y=mx+b
You are told m=1/2 and y=-3 when x=0
-3=(1/2)0+b
b=-3
EQUATION:
y=(1/2)x-3
================
Cheers,
Stan H.